# FixAgent: Hierarchical Multi-Agent Framework for Unified Software Debugging

### v2

@@ Abstract
Software debugging is a time-consuming endeavor involving a series of steps, such as fault localization and patch generation, each requiring thorough analysis and a deep understanding of the underlying logic. While large language models (LLMs) demonstrate promising potential in coding tasks, their performance in debugging remains limited. Current LLM-based methods often focus on isolated steps and struggle with complex bugs. In this paper, we propose the first end-to-end framework, FixAgent, for unified debugging through multi-agent synergy. It mimics the entire cognitive processes of developers, with each agent specialized as a particular component of this process rather than mirroring the actions of an independent expert as in previous multi-agent systems. Agents are coordinated through a three-level design, following a cognitive model of debugging, allowing adaptive handling of bugs with varying complexities. Experiments on extensive benchmarks demonstrate that FixAgent significantly outperforms state-of-the-art repair methods, fixing 1.25$\times$ to 2.56$\times$ bugs on the repo-level benchmark, Defects4J. This performance is achieved without requiring ground-truth root-cause code statements, unlike the baselines. Our source code is available on an anonymous link: https://github.com/AcceptePapier/UniDebugger.
FixAgent: Hierarchical Multi-Agent Framework for Unified Software Debugging
Cheryl Lee1, Chunqiu Steven Xia2, Longji Yang1, Jen-tse Huang1, Zhouruixing Zhu3, Lingming Zhang2, Michael R. Lyu1 1The Chinese University of Hong Kong, 2University of Illinois Urbana-Champaign, 3The Chinese University of Hong Kong, Shenzhen cheryllee@link.cuhk.edu.hk, chunqiu2@illinois.edu, 1155191588@link.cuhk.edu.hk, jthuang@cse.cuhk.edu.hk, zhouruixingzhu@link.cuhk.edu.cn, lingming@illinois.edu, lyu@cse.cuhk.edu.hk
@@ 1 Introduction
Debugging is a crucial process of identifying, analyzing, and rectifying bugs or issues in software. Significant advancements (Yang et al., 2024; Xia and Zhang, 2022; Wei et al., 2023; Xia and Zhang, 2024) have been achieved in addressing bugs with the boost of LLMs—they typically propose a prompting framework and query an LLM to automate an isolated phase in debugging, generally Fault Localization (FL) or Automated Program Repair (APR). FL attempts to identify suspicious code statements, while APR provides patches or fixed code snippets. A typical “diff” patch records the textual differences between two source code files, as shown in Figure 1. The goal of debugging is to apply such a patch to fix the bug accurately.
Figure 1: An example of the “diff” patch, where “@@ -142,5 +142 5 @@” indicates line 142 is modified.
@@
Figure 2: Overview of FixAgent. It starts with the simple L1 repair. If no plausible patch is generated, the L2 repair is triggered, and so is L3. Agents on the same level can communicate with others.
While LLMs demonstrate great potential in addressing individual debugging tasks for basic bugs, previous studies cannot offer a satisfying solution for the entire debugging task. To bridge this gap, we propose the first multi-agent framework, FixAgent, for end-to-end unified debugging. A key challenge lies in how to coordinate multiple agents. Previous studies that employ LLM-based multi-agent frameworks for complex problem-solving typically regard each agent as an expert and imitate the collaborative practice in humans (Hong et al., 2024; Islam et al., 2024; Qian et al., 2024). However, such a paradigm has limitations in debugging. It overly emphasizes mimicking human collaboration, thus not suitable for the logical, stepwise nature of debugging. Furthermore, it faces challenges in the fine-grained allocation and optimization of resources regarding problem difficulties. The communication links among agents are mesh-like and extensive, leading to unnecessary high communication redundancy.
@@
In response, we uniquely structure FixAgent as a hierarchical multi-agent coordination paradigm, as shown in Figure 2. Our key insight is that an agent should act as a specific component of the cognitive process of debugging rather than mirroring the actions of an independent expert in teamwork. In teamwork, members bring diverse perspectives and backgrounds. They engage in bidirectional discussion and negotiation to share knowledge and solve tasks collaboratively. In contrast, the cognitive process is unidirectional and coherent, with each stage building on the last. It accumulates knowledge and employs tools toward a common objective, primarily relying on self-reflection to guide thought. Since the latter aligns seamlessly with debugging, we ground the multi-agent coordination in Hale and Haworth’s cognitive debugging model (Hale and Haworth, 1991). It argues that developers employ a multi-level goal-orientated mechanism during debugging based on structural learning theory (Hale et al., 1999). The initial level only provides quick and straightforward solutions, and if that fails, higher levels of repairs are triggered for complex bugs, entailing more cognitive activities that involve deeper analysis, tool invocation, and external information ingestion.
@@
Particularly, FixAgent first isolates the code file most likely causing the error through unit testing. On level one (L1), only Locator and Fixer are initiated for simple bugs. If the generated patch is not plausible (i.e., pass all test cases), Summarizer and Slicer are triggered with the bug-located code file on level two (L2). Then Locator, Fixer, and FixerPro are called sequentially with access to the responses from Summarizer and Slicer. If it still fails, FixAgent undertakes a deeper inspection and thus activates all agents (i.e., level three, L3), where Helper searches for solutions online to guide all the other agents, and RepoFocus examines the entire program to provide a list of bug-related code files, followed by the other agents working in line. Slicer, Locator, and Fixer can invoke traditional code analysis tools on the last two levels to collect static and dynamic analysis information.
We evaluate FixAgent on four benchmarks with bug-fix pairs across three programming languages, including real-world software (Defects4J Just et al., 2014) and competition programs (QuixBugs Lin et al., 2017, Codeflaws Tan et al., 2017, ConDefects Wu et al., 2024). On Defects4J, FixAgent achieves a new state-of-the-art (SoTA) by correctly fixing 197 bugs with 286 plausible fixes. Our method does not require prior FL while still outperforming strong baselines like ChatRepair, which uses ground-truth root causes with 10x sampling times. Plus, FixAgent fixes all bugs in QuixBugs and generates 2.2$\times$ more plausible fixes on Codeflaws. Our ablation study shows that FixAgent consistently makes significant improvements across various LLM backbones.
@@ 2 Related Work
Software Debugging: FL and APR are the two core steps in software debugging Benton et al. (2020). Many studies aim to automate either step. Traditional FL are typically spectrum-based (Abreu et al., 2009, 2006; Zhang et al., 2011) or mutation-based (Moon et al., 2014; Li and Zhang, 2017; Zhang et al., 2013). Learning-based FL learns program behaviors from rich data sources (Li et al., 2021, 2022; Lou et al., 2021; Li et al., 2019) via multiple types of neural networks. Recently, LLMAO (Yang et al., 2024) proposes to utilize LLMs for test-free FL. APR research either searches for suitable solutions from possible patches (Goues et al., 2012; Weimer et al., 2009a, b) or directly generates patches by representing the generation as an explicit specification inference (Mechtaev et al., 2016; Nguyen et al., 2013; Liu et al., 2019). Learning-based studies ((Lutellier et al., 2020; Jiang et al., 2021; Ye et al., 2022)) translate faulty code to correct code using neural machine translation (NMT). On top of directly prompting LLMs (Xia et al., 2023), recent studies have explored prompt engineering (Xia and Zhang, 2024, 2022) or combining code synthesis (Wei et al., 2023) to improve LLM for APR.
Large Language Models: Various LLMs have been developed for code synthesis (Chen et al., 2021; Guo et al., 2024; Rozière et al., 2023), in addition to general-purpose LLMs (OpenAI, 2023a, b; Anil et al., 2023). They have shown potential in solving coding-related tasks, including program repair, as investigated in empirical studies (Xia et al., 2023; Huang et al., 2023; Tian et al., 2024).
@@
LLM-based Multi-Agent (LLM-MA): The inspiring capabilities of the single LLM-based agent boost the development of multi-agent frameworks. Recent research has demonstrated promising results in utilizing LLM-MA for complex problem-solving, including software engineering, science, and other society-simulating activities. For software engineering, LLM-MA studies (Qian et al., 2024; Hong et al., 2024; Dong et al., 2023; Huang et al., 2024; Islam et al., 2024) usually emulate real-world roles (e.g., product managers, programmers, and testers) and collaborate through communication. Hong et al. (2024); Dong et al. (2023) adopt a shared information pool to reduce overhead.
3 FixAgent
@@
FixAgent is an end-to-end framework for unified debugging through LLM-based multi-agent synergy. It comprises seven agents, each specialized as a state in Hale and Haworth’s cognitive debugging model with an explicit goal instead of being a task-oriented, individual entity. FixAgent operates on a three-level architecture, initializing different levels of repair involving different agents, and agents can adaptively invoke tools in a given toolbox. The communication among agents on the same level follows an assembly-line thought process rather than the mesh interactions typical of teams. This section introduces the profiles of agents, their interactions with external environments, and the hierarchical organization. Prompt details are displayed in A.3.
@@ 3.1 Profiles of Agents
All agents in FixAgent share a one-shot structure of the system prompt, which defines their roles, skills, actions, objectives, and constraints, followed by a manually crafted example to illustrate the desired response format, as shown in Figure 3. Moreover, all agents have access to certain meta-information known as bug metadata, including the bug-located code file, failing test cases, reported errors during compilation and testing, and program requirements described in natural language. The response of each agent comprises two elements: the answer fulfilling its goal and an explanation of its thinking. Inspired by a software engineering principle, rubber duck debugging (Andrew and David, 2000), where developers articulate their expectations versus actual implementations to identify the gap, we request each agent to monitor key program variables and explain how it guides the answer.
Figure 3: A one-shot system prompt defines the role, skills, actions, objective, and constraints of the agent.
Helper. The goal of Helper is to provide references via retrieve-augmented generation (RAG), inspired by the fact that developers commonly utilize web search engines (e.g., Google) to enhance productivity (Xia et al., 2017). By analyzing the bug metadata, Helper generates a short query and invokes an external search engine to retrieve the best-matching solutions. Then, it integrates the retrieval results and generates reference solutions that are customized for the context of the buggy code (e.g., variable naming and function structure). These are internal processes hidden from the other agents, and only the final response—a reference solution—can be read only on L3. Figure 4 provides an example of the response from Helper.
@@
Fixer. Fixer is prompted to generate a ‘diff‘, as shown in Figure 1 in Introduction. It receives code marked by Locator and other bug metadata. To maintain consistency between Locator and Fixer, Fixer first assesses whether the marked lines should be modified and then describes the modification. This also enables Fixer to correct possible errors made by Locator. On L2 and L3, Fixer also has access to static/dynamic analysis and auxiliary information generated by upstream agents.
@@ 3.2 External Interactions
Plausibility Feedback. We conduct repairs by modifying the buggy code rather than directly applying the generated patches because they usually make mistakes in indices of code lines, as LLMs often struggle with counting, and the patch may be generated based on a sliced segment rather than the original code, causing changes in the indices of code lines. Our modification is rule-based by matching the invariant context between the patch and the buggy code. Afterward, we compile and run tests on the fixed program. The patch will be sent to developers for manual correctness verification if passing all test cases. The results of plausibility testing are returned to FixAgent as feedback for further processing, such as initializing a higher level of repair or conducting self-reflection.
Tool Usage. The toolbox of FixAgent currently contains static analysis tools, dynamic analysis tools, and a search engine optimized for LLMs and RAG. Static analysis provides static errors and warnings, as well as Abstract Syntax Tree (AST), a tree representation of the abstract syntactic structure of source code. We mainly consider the coverage of failing test cases in dynamic analysis since it is commonly assumed that code statements not executed during testing are less likely to cause the failures Kochhar et al. (2016). Note that the results of tool invocations will be stored for the requests of downstream agents to avoid duplicated invocations.
@@ 3.3 Hierarchical Coordination
The first level identifies and fixes obvious logic errors in code, so we assume that agents can easily fix the bug. Thus, L1 only contains Locator and Fixer with a simple communication flow, where Fixer generates a patch based on the fault localization done by Locator. If the fix is not plausible, then L2 is triggered, assuming that the fix can be achieved by only inspecting a single file, which is often applied in previous FL and APR studies (Soremekun et al., 2023). L2 involves five agents: Slicer, Summarizer, Locator, Fixer, and FixerPro. Slicer isolates a suspicious segment, and Summarizer generates a code summary simultaneously, both from the bug-located file. Afterward, Locator, Fixer, and FixerPro work in sequence based on the suspicious segment, and all have access to the code summary. If it still fails, we turn to L3, which believes that the bug is very complex, so repairing it requires understanding cross-file dependencies and external information. L3 activates all the seven agents. First, Helper and RepoFocus are initialized at the same time, and Helper provides references for all other agents. Then, Summarizer runs multiple times following the file list identified by RepoFocus. These code summaries are shared by the remaining four agents, which work one after another subsequently with the same workflow as L2. Figure 5 presents an example conversation among agents on L3.
Upon no plausible patch, FixAgent gradually requests agents on L3 to reflect and refine its answer based on the plausibility feedback in a reversed order, as downstream agents are more error-prone as errors may accumulate during the debugging process. Specifically, we first ask FixerPro to refine its response. If the fix fails again, we then request Fixer, leading to input changes of FixerPro, so it is re-sampled the second time. Eventually, plausible patches are returned to developers for manual verification. If no plausible patch is produced, FixAgent will display the analysis report written by FixerPro and take another run until reaching a resource threshold.
@@
Tools Sampling Times Defects4J-Java Codeflaws-C QuixBugs-Java QuixBugs-Python Note #Corr #Plau #Corr #Plau #Corr #Plau #Corr #Plau Angelix 1,000 - - 318 591 - - - - Realistic FL Prophet 1,000 - - 310 839 - - - - SPR 1,000 - - 283 783 - - - - CVC4 - - - 15 91 - - - - Semfix 1,000 25 68 38 56 - - - - Recoder 100 72 140 - - 17 17 - - GenProg 1000 5 20 255-369 1423 1 4 - - Perfect FL CoCoNuT 20,000 44 85 423 716 13 20 19 21 CURE - 57 104 - - 26 35 - - RewardRepair 200 90 75 - - 20 - - - Tbar 500 77 121 - - - - - - AlphaRepair 5,000 110 159 - - 28 30 27 32 Repilot 5,000 116 - - - - - - - ChatRepair 100-200 157 - - - 40 40 40 40 CodeLlama-34b 20 24 41 91(%) 1,488 25 28 33 33 Perfect FL LLaMA2-70b 39 78 91(%) 1,576 25 28 33 33 DeepSeekCoder 57 82 93(%) 1,937 30 34 25 38 gemini-1.5-flash 19 36 86(%) 1,291 29 32 29 35 gpt-3.5-turbo-ca 45 71 94(%) 2,343 33 34 34 36 claude-3.5-sonnet 70 116 95(%) 2,624 36 37 40 40 gpt-4o 72 119 93(%) 2,549 35 36 39 39 FixAgent-Lite 5 - - 95(%) 3,130 40 40 40 40 FixAgent 20 197 286 - - - - - -

### v3

# UniDebugger: Hierarchical Multi-Agent Framework for Unified Software Debugging

@@ Abstract
[…]
UniDebugger: Hierarchical Multi-Agent Framework for Unified Software Debugging
Cheryl Lee1*, Chunqiu Steven Xia2, Longji Yang1, Jen-tse Huang1, Zhouruixing Zhu3, Lingming Zhang2, Michael R. Lyu1 1The Chinese University of Hong Kong    2University of Illinois Urbana-Champaign 3The Chinese University of Hong Kong, Shenzhen 1*cheryllee@link.cuhk.edu.hk  2chunqiu2@illinois.edu   6lingming@illinois.edu   7lyu@cse.cuhk.edu.hk
@@ 1 Introduction
Debugging is a crucial process of identifying, analyzing, and rectifying bugs in software. Significant advancements (LLMAO; AlphaRepair; Repilot; ChatRepair) have been achieved in addressing bugs with the boost of LLMs—they typically propose a prompting framework and query an LLM to automate an isolated phase in test-driven debugging, generally Fault Localization (FL) or Automated Program Repair (APR). FL attempts to identify suspicious code statements, while APR provides patches or fixed code snippets. A typical “diff” patch records the textual differences between two source code files, as shown in Figure 1.
Figure 1: An example of the “diff” patch.
@@
Figure 2: Overview of UniDebugger. It starts with the simple L1 repair. If no plausible patch is generated, the L2 repair is triggered, and so is L3. Agents on the same level can communicate with others.
While LLMs demonstrate great potential in addressing individual debugging tasks for basic bugs, previous APR studies cannot deliver a satisfying end-to-end solution for the entire task. To bridge this gap, we propose the first multi-agent framework, UniDebugger, for developer-side debugging, which can be seamlessly integrated into the CI/CD pipeline. A key challenge lies in how to coordinate multiple agents. Existing multi-agent frameworks for complex problem-solving adopt a horizontal collaboration paradigm where each agent acts as an independent expert, mirroring human team dynamics through bidirectional discussions and mesh-like communication (e.g.,  MetaGPT; MapCoder; ChatDev). However, this design introduces critical limitations for debugging: (1) The inherent redundancy of peer-to-peer negotiations conflicts with debugging’s logical, incremental nature, wasting computational resources on trivial bugs; (2) The lack of structured problem decomposition leads to suboptimal resource allocation across bugs of varying complexity.
@@
In response, we uniquely structure UniDebugger as a hierarchical multi-agent coordination paradigm grounded in cognitive debugging theory, as shown in Figure 2. Our key innovation lies in redefining agents as functional components of a unified cognitive process rather than autonomous experts. Inspired by Hale and Haworth’s model of structural learning (Hale), which argues that developers employ a multi-level goal-orientated mechanism during debugging (HaleEvaluation), UniDebugger implements a three-level mechanism. The initial level only provides quick and straightforward solutions, and if that fails, higher levels of repairs are triggered for complex bugs, entailing more cognitive activities that involve deeper analysis, tool invocation, and external information ingestion. Unlike mesh-based teamwork, where members with diverse perspectives and backgrounds engage in bidirectional discussion and negotiation to share knowledge and solve tasks collaboratively, this cognitive process is unidirectional and coherent, with each stage accumulating knowledge incrementally and building on the last. By seamlessly aligning with debugging, this paradigm serves as the basis of our design for multi-agent synergy.
@@
Particularly, UniDebugger first isolates the code file most likely causing the error through unit testing. On level one (L1), only Locator and Fixer are initiated for simple bugs. If the generated patch is not plausible (i.e., pass all test cases), Summarizer and Slicer are triggered with the bug-located code file on level two (L2). Then Locator, Fixer, and FixerPro are called sequentially with access to the responses from Summarizer and Slicer. If it still fails, UniDebugger undertakes a deeper inspection and thus activates all agents (i.e., level three, L3), where Helper searches for solutions online to guide all the other agents, and RepoFocus examines the entire program to provide a list of bug-related code files, followed by the other agents working in line. Slicer, Locator, and Fixer can invoke traditional code analysis tools on the last two levels to collect static and dynamic analysis information.
[…]
@@ 2 Related Work
Software Debugging: Test-driven debugging has gained popularity since testing is an essential and prevalent part of practical CI/CD, and this methodology typically contains two core steps UnifiedSurvey: FL and APR. Many studies aim to automate either step. Traditional FL are typically spectrum-based (SpectrumFL; Ochiai; SpectrumFL2011) or mutation-based (MutantFL; TraPT; InjectFL). Learning-based FL learns program behaviors from rich data sources (DeepRL4FL; FixLocator; GRACE; DeepFL) via multiple types of neural networks. Recently, LLMAO (LLMAO) proposes to utilize LLMs for test-free FL. APR research either searches for suitable solutions from possible patches (genProg1; genProg2; geneticAPR) or directly generates patches by representing the generation as an explicit specification inference (Angelix; SemFix; Tbar). Learning-based studies ((CoCoNuT; CURE; RewardRepair)) translate faulty code to correct code using neural machine translation (NMT). On top of directly prompting LLMs (LLMAPREra), recent studies have explored prompt engineering (ChatRepair; AlphaRepair) or combining code synthesis (Repilot) for APR. In addition to developer-oriented debugging, techniques for addressing user-oriented issues receive great attention Agentless; SWE-Bench. These methods responsively update software based on problems discovered by users. This paper focuses on the CI phase, reducing the burden on developers responsible for proactively detecting and resolving bugs to prevent errors from entering production.
@@
LLM-based Multi-Agent (LLM-MA): Various LLMs have been developed for code synthesis (Codex; DeepSeekCoder; CodeLlama), in addition to general-purpose LLMs (GPT3-5; GPT4; Gemini). They have shown potential in solving coding-related tasks, including program repair (LLMAPREra; LLMCSurvey; DebugBench). The inspiring capabilities of the single LLM-based agent boost the development of multi-agent frameworks. Recent research has demonstrated promising results in utilizing LLM-MA for complex problem-solving, including software engineering, science, and other society-simulating activities. For software engineering, LLM-MA studies (ChatDev; MetaGPT; Self-collaboration; AgentCoder; MapCoder) usually emulate real-world roles (e.g., product managers, programmers, and testers) and collaborate through communication. MetaGPT; Self-collaboration adopt a shared information pool to reduce overhead.
3 UniDebugger
@@
UniDebugger is an end-to-end framework for unified debugging through LLM-based multi-agent synergy. It comprises seven agents, each specialized as a state in Hale and Haworth’s cognitive debugging model with an explicit goal instead of being a task-oriented, individual entity. UniDebugger operates on a three-level architecture, initializing different levels of repair involving different agents, and agents can adaptively invoke tools in a given toolbox. The communication among agents on the same level follows an assembly-line thought process rather than the mesh interactions typical of teams. This section introduces the profiles of agents, their interactions with external environments, and the hierarchical organization. Prompt details are displayed in A.3.
@@ 3.1 Profiles of Agents
All agents in UniDebugger share a one-shot structure of the system prompt, which defines their roles, skills, actions, objectives, and constraints, followed by a manually crafted example to illustrate the desired response format. Moreover, all agents have access to certain meta-information known as bug metadata, including the bug-located code file, failing test cases, reported errors during compilation and testing, and program requirements described in natural language. The response of each agent comprises two elements: the answer fulfilling its goal and an explanation of its thinking. Inspired by a software engineering principle, rubber duck debugging (RubberDuck), where developers articulate their expectations versus actual implementations to identify the gap, we request each agent to monitor key program variables and explain how it guides the answer.
Helper. The goal of Helper is to provide references via retrieve-augmented generation (RAG), inspired by the fact that developers commonly utilize web search engines (e.g., Google) to enhance productivity (DevSearch). By analyzing the bug metadata, Helper generates a short query and invokes an external search engine to retrieve the best-matching solutions. Then, it integrates the retrieval results and generates reference solutions that are customized for the context of the buggy code (e.g., variable naming and function structure). These are internal processes hidden from the other agents, and only the final response—a reference solution—can be read only on L3. Figure 3 provides an example of the response from Helper.
@@
Fixer. Fixer is prompted to generate a "diff", as shown in Figure 1 in Introduction. It receives code marked by Locator and other bug metadata. To maintain consistency between Locator and Fixer, Fixer first assesses whether the marked lines should be modified and then describes the modification. This also enables Fixer to correct possible errors made by Locator. On L2 and L3, Fixer also has access to static/dynamic analysis and auxiliary information generated by upstream agents.
@@ 3.2 External Interactions
Plausibility Feedback. We conduct repairs by modifying the buggy code rather than directly applying the generated patches because they usually make mistakes in the format, especially the indices of code lines, as LLMs often struggle with counting. Unlike prior studies, such as  Agentless, using a Search/Replace method for code editing, our post-adjustment methodology more readily aligns with LLM training databases as the outcomes of extensive code edits are stored in a "diff" format. Our modification is rule-based by matching the invariant context between the patch and the buggy code. Afterward, we compile and run tests on the fixed program. The patch will be sent to developers for manual correctness verification if passing all test cases. The results of plausibility testing are returned to UniDebugger as feedback for further processing, such as initializing a higher level of repair or conducting self-reflection.
Tool Usage. The toolbox of UniDebugger currently contains static analysis tools, dynamic analysis tools, and a search engine optimized for LLMs and RAG. Static analysis provides static errors and warnings, as well as Abstract Syntax Tree (AST), a tree representation of the abstract syntactic structure of source code. We mainly consider the coverage of failing test cases in dynamic analysis since it is commonly assumed that code statements not executed during testing are less likely to cause the failures FLSurvey. Note that the results of tool invocations will be stored for the requests of downstream agents to avoid duplicated invocations.
@@ 3.3 Hierarchical Coordination
The first level identifies and fixes obvious logic errors in code, so we assume that agents can easily fix the bug. Thus, L1 only contains Locator and Fixer with a simple communication flow, where Fixer generates a patch based on the fault localization done by Locator. If the fix is not plausible, then L2 is triggered, assuming that the fix can be achieved by only inspecting a single file, which is often applied in previous FL and APR studies (Assumptions). L2 involves five agents: Slicer, Summarizer, Locator, Fixer, and FixerPro. Slicer isolates a suspicious segment, and Summarizer generates a code summary simultaneously, both from the bug-located file. Afterward, Locator, Fixer, and FixerPro work in sequence based on the suspicious segment, and all have access to the code summary. If it still fails, we turn to L3, which believes that the bug is very complex, so repairing it requires understanding cross-file dependencies and external information. L3 activates all the seven agents. First, Helper and RepoFocus are initialized at the same time, and Helper provides references for all other agents. Then, Summarizer runs multiple times following the file list identified by RepoFocus. These code summaries are shared by the remaining four agents, which work one after another subsequently with the same workflow as L2. Figure 4 presents an example conversation among agents on L3.
Upon no plausible patch, UniDebugger gradually requests agents on L3 to reflect and refine its answer based on the plausibility feedback in a reversed order, as downstream agents are more error-prone as errors may accumulate during the debugging process. Specifically, we first ask FixerPro to refine its response. If the fix fails again, we then request Fixer, leading to input changes of FixerPro, so it is re-sampled the second time. Eventually, plausible patches are returned to developers for manual verification. If no plausible patch is produced, UniDebugger will display the analysis report written by FixerPro and take another run until reaching a resource threshold.
@@
Tools Sampling Times Defects4J-Java Codeflaws-C QuixBugs-Java QuixBugs-Python Note #Corr #Plau #Corr #Plau #Corr #Plau #Corr #Plau Angelix 1,000 - - 318 591 - - - - Realistic FL Prophet 1,000 - - 310 839 - - - - SPR 1,000 - - 283 783 - - - - CVC4 - - - 15 91 - - - - Semfix 1,000 25 68 38 56 - - - - Recoder 100 72 140 - - 17 17 - - GenProg 1000 5 20 255-369 1423 1 4 - - Perfect FL CoCoNuT 20,000 44 85 423 716 13 20 19 21 CURE - 57 104 - - 26 35 - - RewardRepair 200 90 75 - - 20 - - - Tbar 500 77 121 - - - - - - AlphaRepair 5,000 110 159 - - 28 30 27 32 Repilot 5,000 116 - - - - - - - ChatRepair 100-200 157 - - - 40 40 40 40 CodeLlama-34b 20 24 41 91(%) 1,488 25 28 33 33 Perfect FL LLaMA2-70b 39 78 91(%) 1,576 25 28 33 33 DeepSeekCoderV2 57 82 93(%) 1,937 30 34 25 38 gemini-1.5-flash 19 36 86(%) 1,291 29 32 29 35 gpt-3.5-turbo-ca 45 71 94(%) 2,343 33 34 34 36 claude-3.5-sonnet 70 116 95(%) 2,624 36 37 40 40 gpt-4o 72 119 93(%) 2,549 35 36 39 39 UniDebugger-Lite 5 - - 95(%) 3,130 40 40 40 40 UniDebugger 20 197 286 - - - - - -